# Dynamic Rank Adjustment in Diffusion Policies for Efficient and Flexible Training

### v2

@@ VIII-A Limitations
Additionally, while we have tested and evaluated various decay functions for the rank scheduler, the current implementation of dynamic rank adjustment in the DRIFT framework follows a monotonic schedule. Although we have conducted ablation studies on decay functions and terminal ranks, the impact of these design choices is likely task-dependent. Furthermore, the rank adjustment of different convolutional blocks in this work is applied uniformly throughout the U-Net backbone, even though different blocks may have varying highest possible ranks. Future research could explore more adaptive and intelligent strategies for adjusting trainable ranks to enhance training efficiency and performance, as well as identify suitable decay functions and terminal ranks for scenarios beyond those covered in this work.

### v3

@@ VIII-A Limitations
Additionally, while we have tested and evaluated various decay functions for the rank scheduler, the current implementation of dynamic rank adjustment in the DRIFT framework follows a monotonic schedule. Although we have conducted ablation studies on decay functions and terminal ranks, the impact of these design choices is likely task-dependent.
Furthermore, the rank adjustment of different convolutional blocks in this work is applied uniformly throughout the U-Net backbone, even though different blocks may have varying highest possible ranks. Future research could explore more adaptive and intelligent strategies for adjusting trainable ranks to enhance training efficiency and performance, as well as identify suitable decay functions and terminal ranks for scenarios beyond those covered in this work.
@@ VIII-B Implications
Acknowledgments
This work was supported by Office of Naval Research award N00014-24-1-2124